# Rewriting Pre-Training Data Boosts LLM Performance in Math and Code

### v1

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable zero-shot and few-shot capabilities across diverse tasks, yet their proficiency in mathematical reasoning and program synthesis remains constrained by the quality of pre-training corpora. Existing public datasets for specialized domains, such as The-Stack-v1 and v2 for code [1, 2] and Finemath-4+ for mathematics [3], rely primarily on rule-based extraction from web crawls (e.g., CommonCrawl) [4] or model-based scoring to filter low-quality samples. However, these approaches often retain noisy, redundant, or stylistically inconsistent data, limiting their effectiveness. For instance, as shown in Figure 1, continual pre-training of Llama-3.1-8B [5] on The-Stack-v1/v2 typically maintains baseline performance on benchmarks like HumanEval and HumanEval+ but struggles to achieve significant gains due to unaddressed data quality issues. Unlike prior methods that focus on filtering or preserving original samples, we propose rewriting pre-training corpora to eliminate noise and redundancy, delivering high-quality, self-contained data that enables efficient model learning.
@@
This paper introduces two openly licensed datasets under the Llama 3.3 Community License, designed to advance LLM capabilities in code generation and mathematical reasoning. SwallowCode ($\approx$16.1 billion tokens) refines Python snippets from The-Stack-v2 through a novel four-stage pipeline: sequential syntax validation, pylint-based style filtering, and a two-stage LLM rewriting process that enforces style conformity and transforms snippets into algorithmically efficient, self-contained examples. By rewriting, rather than merely filtering, we eliminate noise and redundancy that persist in datasets like Stack-Edu, providing high-quality data that enhances model learning efficiency (Figure 1). SwallowMath ($\approx$2.3 billion tokens) transforms Finemath-4+ by removing boilerplate, restoring missing context, and reformatting solutions into concise and step-by-step explanations. Although our experiments focus on Python for ease of automated evaluation, the pipeline is language-agnostic: any programming language with a parseable syntax and a linter can benefit from the same treatment, broadening its applicability to diverse coding domains.
@@
Continual pre-training of Llama-3.1-8B [5] for 50 billion tokens with SwallowCode and multilingual datasets improves pass@1 by +17.0 on HumanEval and +17.7 on HumanEval+ compared to an equivalent budget using Stack-Edu [3]. Similarly, substituting Finemath-4+ with SwallowMath yields +12.4 accuracy on GSM8K and +7.6 on MATH. To ensure the robustness of these gains, we conducted rigorous checks for test-set leakage, streaming the entire 16.1B-token SwallowCode corpus and finding no exact matches or high-similarity documents (Jaccard similarity $\geq$ 0.8) with HumanEval or HumanEval+ prompts. All datasets, prompts, and checkpoints are publicly released, enabling reproducible research. SwallowCode and SwallowMath provide a scalable framework for enhancing LLM pre-training, driving advancements in automated reasoning and software development.
@@ 2.1 Open code corpora
The Stack v1 [1] aggregates approximately 3TB of permissively licensed source code from public GitHub repositories, spanning 358 programming languages. While deduplicated and vetted for licensing, its filtering is primarily syntactic, retaining boilerplate and auto-generated files without language-specific refinement. This limits its utility for training models on high-quality, semantically rich code. The Stack v2 [2] extends this pipeline by sourcing data from the Software Heritage archive, relaxing licensing constraints, and introducing language-specific filters to enhance data quality. With a scale of approximately 900 billion tokens, it represents a significant increase in volume over v1. However, like its predecessor, v2 adopts a selection-only approach, preserving files without semantic enhancements, which leaves inconsistent coding styles and fragmented scripts unaddressed.
@@ 2.2 Classifier-based filtering for code corpora
Recent work has shown that classifier-based filtering strategies, such as the FineWeb-Edu approach [6], can be effective for curating high-quality datasets [7]. These methods aim to improve model performance on code-related tasks by selecting semantically rich and well-documented samples from large-scale corpora. In this context, Stack-Edu [3] represents a significant effort to create a filtered variant of StarCoder2Data [2] prioritizing high-quality code.
Stack-Edu begins by selecting the 15 most prevalent programming languages from StarCoder2Data, forming a subset of approximately 450 billion tokens. To assess code quality, Stack-Edu leverages Llama-3-70B-Instruct [5] to generate synthetic annotations for 500,000 code fragments, rating each on a 0–5 scale based on educational and structural quality. These annotations train language-specific classifiers based on the StarEncoder model [8], achieving F1 scores above 0.7 for most languages when applying a threshold of 3 for binary classification. The resulting dataset comprises 125 billion tokens and demonstrates improved model convergence and higher pass@1 scores on HumanEval compared to unfiltered corpora [3]. However, Stack-Edu’s exclusionary filtering strategy discards low-scoring snippets rather than rewriting or augmenting them, leaving residual issues—such as missing context or inconsistent naming conventions in the retained data.
@@ 2.3 LLM-driven pre-training corpus rewriting
Efforts to enhance code datasets using large language models (LLMs) have gained traction. Austin et al. [9] explored LLM capabilities for program synthesis, finding that human feedback significantly improves code accuracy, but their work focused on synthesis rather than dataset rewriting. Cosmopedia [10] demonstrated the potential of synthetic data generation, using Mixtral-8×7B-Instruct [11] to create high-quality text corpora, though it did not address code-specific challenges.
Jain et al. [12] applied LLM-driven transformations to instruction-tuning exemplars, focusing on variable renaming, modularization, and comment addition. While effective for fine-tuning, their approach is limited in scope and context compared to our SwallowCode pipeline. Their transformations address only a subset of stylistic improvements, whereas our Style-Guided Code Rewriting (SGCR) pipeline comprehensively enforces ten criteria from the Google Python Style Guide, including descriptive variable naming, effective type annotations, modular function design, error handling, and readability-focused formatting (see Section 3.3.1). Furthermore, our Self-Contained Optimization Rewriting (SCOR) pipeline introduces semantic enhancements—ensuring self-containment by resolving dependencies, optimizing algorithms, and transforming trivial snippets into instructive examples—which are absent in Jain et al.’s work (see Section 3.3.2). Additionally, Jain et al. target instruction-tuning data, a smaller and more curated dataset, whereas SwallowCode systematically rewrites large-scale pre-training corpora, a more challenging task due to the diversity and volume of code samples. Our pipeline integrates rigorous preprocessing with syntax error and linter-based filtering to ensure high-quality inputs for rewriting, and evaluates LLM-based scoring in ablation studies, comparing its computational cost and effectiveness against rewriting (see Section 3.2.3). In contrast, Jain et al. neither employ filtering nor assess the cost-effectiveness of their transformations.
SwallowCode’s transform-and-retain paradigm upgrades low-quality code rather than discarding it, addressing limitations of existing datasets like The Stack [1, 2], Stack-Edu [7, 3], and OpenCoder [13], which rely on exclusionary filtering. By combining filtering, SGCR, and SCOR, SwallowCode produces a high-quality corpus that significantly improves performance on HumanEval and HumanEval+ benchmarks, advancing the state-of-the-art in code dataset curation.
@@ 3 Construction of the code corpus
The development of SwallowCode is driven by an empirical approach, informed by data ablation experiments evaluating each stage of the data processing pipeline, as illustrated in Figure 2. These experiments are detailed in Appendix H.1, which presents the detailed results of ablation experiments. In this section, we present the experimental results and describe the design choices shaping the corpus. All ablation experiments follow an iterative dataset construction process, in which the baseline model for each subsection incorporates only the processing steps described in prior subsections, unless explicitly stated otherwise. To ensure full reproducibility, all code, configurations, and supporting materials are publicly available at https://github.com/rioyokotalab/swallow-code-math.
@@ 3.1 Experimental setup
To evaluate the impact of each design choice in our data processing pipeline, we conduct systematic data ablation studies. Each ablation trains a model that differs only in the target pre-training dataset, holding all other factors constant, including model architecture, parameter count, non-target data, total token budget, and hyperparameters. Specifically, we perform continual pre-training starting from Llama-3.1-8B, using a total of approximately 50 billion tokens, a maximum sequence length of 8,192, a global batch size of approximately 4 million tokens, and the Llama-3 tokenizer. The target dataset is processed with less than one epoch within each ablation. We evaluate model checkpoints approximately every 10 billion tokens using a consistent suite of downstream benchmarks.
Continual pre-training is conducted using Megatron-LM (version core_r0.9.0) [14]. For evaluation, we utilize the BigCodeBench [15] and lm-evaluation-harness [16] on the following benchmarks: OpenBookQA [17], TriviaQA [18], HellaSwag [19], SQuAD 2.0 [20], XWINO [21], MMLU [22], GSM8K [23], BBH [24], HumanEval [25], and HumanEval+ [26]. The effectiveness of the code corpus is specifically assessed using HumanEval and HumanEval+.
The pre-training data mixture consists of 84% multilingual text and 16% code. Detailed proportions and data source are provided in Appendix A.4.1, with comprehensive training hyperparameters reported in Appendix A.1. All ablation models, associated checkpoints, and supporting materials are publicly available in our Hugging Face repository to ensure full reproducibility.
@@ 3.2 Filtering
Figure 3: Performance comparison of filtering methods on HumanEval (left) and HumanEval+ (right) benchmarks. The baseline (Stack v2) is compared against datasets processed with syntax error filtering, linter-based filtering, and LLM-based scoring (evaluated in ablation studies but not adopted). Syntax and linter-based filtering enhance code generation performance, while LLM-based scoring provides marginal gains, considering the computational cost.
@@ 3.2.1 Syntax error filtering
Despite prior processing, the-stack-v2-train-smol-ids [2] contains Python code samples with invalid syntax per Python 3.10 standards. To address this, we apply syntax error filtering by compiling each code sample using Python’s built-in compile() function, discarding any samples that fail to compile. This process reduces the dataset from approximately 41 million to 37 million samples, a 10.6% reduction. As shown in Figure 3, removing syntactically invalid samples improves performance on HumanEval and HumanEval+  [^1] The performance trajectory of The Stack v2 exhibits an initial decline followed by a recovery, consistent with forgetting and subsequent adaptation observed in prior continual pretraining studies [27]. This behavior is not particularly noteworthy.. Consequently, we adopt syntax-error filtering as a standard step in all subsequent experiments.
@@ 3.2.3 LLM-based scoring filtering
Given these limited gains, we compare LLM-based scoring to our LLM-driven rewriting pipeline, which refines code snippets by enhancing clarity and correctness (Section 3.3). Comparing Figure 3 and Figure 4, although the rewriting pipeline requires 1.22 times the computational resources of LLM-based scoring (a 22% increase), it achieves significantly greater performance gains on HumanEval and HumanEval+ (detailed in Appendix F). Consequently, we do not incorporate LLM-based scoring in subsequent experiments, which favors the more effective and comparably efficient rewriting approach.
@@ 3.3 LLM-driven rewriting
Figure 4: Performance comparison of LLM-driven rewriting pipelines on HumanEval (left) and HumanEval+ (right) benchmarks. The baseline (syntax-error and linter-based filtering) is compared against Style-Guided Code Rewriting (SGCR) and Self-Contained Optimization Rewriting (SCOR). SGCR improves performance by over 9 points, while SCOR, applied after SGCR, further enhances scores by over 5 points, demonstrating the effectiveness of stylistic and semantic rewrites in the SwallowCode pipeline.
Recent studies highlight the potential of large language models (LLMs) to transform training data, enhancing the performance of instruction-tuned models. Jain et al. [12] proposed three transformations—variable renaming, modularization, and plan annotation—for code cleaning, demonstrating their effectiveness for instruction tuning. However, their approach is limited to curated instruction-tuning data and a narrow set of stylistic improvements, lacking semantic optimizations or preprocessing. Building on this foundation, our SwallowCode pipeline systematically rewrites large-scale pre-training corpora.
We introduce two complementary LLM-driven rewriting pipelines, Style-Guided Code Rewriting (SGCR) and Self-Contained Optimization Rewriting (SCOR), both adopted in the SwallowCode building pipeline to enhance code quality for pre-training. SGCR prompts an LLM to revise Python code snippets based on criteria from the Google Python Style Guide  [^2] https://google.github.io/styleguide/pyguide.html, enforcing stylistic improvements like clear naming and modular design (Section 3.3.1). SCOR extends SGCR by ensuring self-containment and applying semantic optimizations, such as efficient algorithms and instructive examples (Section 3.3.2). To illustrate the scope of these transformations, Table 1 compares the coverage of SGCR and SCOR against the approach of Jain et al. [12]. While SGCR addresses stylistic criteria like type hints, error handling, and docstrings, SCOR introduces semantic enhancements, including self-containment and optimization(algorithm and data structure), broadening the scope of the SwallowCode pipeline.
Table 1: Comparison of code transformation coverage for Jain et al. [12], Style-Guided Code Rewriting (SGCR), and Self-Contained Optimization Rewriting (SCOR). SGCR addresses stylistic criteria from the Google Python Style Guide, while SCOR extends to semantic optimizations, together forming the comprehensive SwallowCode pipeline.

### v2

@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable zero-shot and few-shot capabilities across diverse tasks, yet their proficiency in mathematical reasoning and program synthesis remains constrained by the quality of pre-training corpora. Existing public datasets for specialized domains, such as The-Stack-v1 and v2 for code [1, 2] and Finemath-4+ for mathematics [3], rely primarily on rule-based extraction from web crawls (e.g., CommonCrawl) [4] or model-based scoring to filter low-quality samples. However, these approaches often retain noisy, redundant, or stylistically inconsistent data, limiting their effectiveness–particularly in the growing trend of multi-stage pretraining (or mid-training) aimed at enhancing mathematical reasoning and program synthesis (e.g., OLMo 2 [5], Nemotron-H [6], and Phi-4 [7]). For instance, as shown in Figure 1, continual pre-training of Llama-3.1-8B [8] on The-Stack-v1/v2 typically maintains baseline performance on benchmarks like HumanEval and HumanEval+ but struggles to achieve significant gains due to unaddressed data quality issues. Unlike prior methods that focus on filtering or preserving original samples, we propose rewriting pre-training corpora to eliminate noise and redundancy, delivering high-quality, self-contained data that enables efficient model learning.
@@
This paper introduces two openly licensed datasets under the Llama 3.3 Community License, designed to advance LLM capabilities in code generation and mathematical reasoning. SwallowCode ($\approx$16.1 billion tokens) refines Python snippets from The-Stack-v2 through a novel four-stage pipeline: sequential syntax validation, pylint-based style filtering, and a two-stage LLM rewriting process that enforces style conformity and transforms snippets into algorithmically efficient, self-contained examples. By rewriting, rather than merely filtering, we eliminate noise and redundancy that persist in datasets like Stack-Edu, providing high-quality data that drives rapid accuracy improvements (Figure 1). SwallowMath ($\approx$2.3 billion tokens) transforms Finemath-4+ by removing boilerplate, restoring missing context, and reformatting solutions into concise and step-by-step explanations. Although our experiments focus on Python for ease of automated evaluation, the pipeline is language-agnostic: any programming language with a parseable syntax and a linter can benefit from the same treatment, broadening its applicability to diverse coding domains.
@@
Continual pre-training of Llama-3.1-8B [8] for 50 billion tokens on mixed dataset with SwallowCode and multilingual corpora improves pass@1 by +17.0 on HumanEval and +17.7 on HumanEval+ compared to an equivalent budget using Stack-Edu [3]. Similarly, substituting Finemath-4+ with SwallowMath yields +12.4 accuracy on GSM8K and +7.6 on MATH. To ensure the robustness of these gains, we conducted rigorous checks for test-set leakage, streaming the entire 16.1B-token SwallowCode corpus and finding no exact matches or high-similarity documents (Jaccard similarity $\geq$ 0.8) with HumanEval or HumanEval+ prompts. All datasets, prompts, and checkpoints are publicly released, enabling reproducible research. SwallowCode and SwallowMath provide a scalable framework for enhancing LLM pre-training, driving advancements in automated reasoning and software development.
@@ 2.1 Open code corpora
The Stack v1 [1] aggregates approximately 3TB of permissively licensed source code from public GitHub repositories. Although deduplication and license filtering are applied, boilerplate and autogenerated files are retained without language-specific refinement. The Stack v2 [2] extends this pipeline by sourcing data from the Software Heritage archive, relaxing licensing constraints, and introducing language-specific filters to enhance data quality. With a scale of approximately 900 billion tokens, it represents a significant increase in volume over v1. However, like its predecessor, v2 adopts a filtering-only approach, preserving files without semantic enhancements, which leaves inconsistent coding styles and fragmented scripts unaddressed.
@@ 2.2 Classifier-based filtering for code corpora
Recent work has shown that classifier-based filtering strategies, such as the FineWeb-Edu approach [9], can be effective for curating high-quality datasets [10]. These methods aim to improve model performance on code-related tasks by selecting semantically rich and well-documented samples from large-scale corpora. In this context, Stack-Edu [3] represents a significant effort to create a filtered variant of StarCoder2Data [2] prioritizing high-quality code. Stack-Edu begins by selecting the 15 most prevalent programming languages from StarCoder2Data, forming a subset of approximately 450 billion tokens. To assess code quality, Stack-Edu leverages Llama-3-70B-Instruct [8] to generate synthetic annotations for 500,000 code fragments, rating each on a 0–5 scale based on educational and structural quality. These annotations train language-specific classifiers based on the StarEncoder model [11], achieving F1 scores above 0.7 for most languages when applying a threshold of 3 for binary classification. The resulting dataset comprises 125 billion tokens and demonstrates improved model convergence and higher pass@1 scores on HumanEval compared to unfiltered corpora [3]. However, Stack-Edu’s exclusionary filtering strategy discards low-scoring snippets rather than rewriting or augmenting them, leaving residual issues—such as missing context or inconsistent naming conventions in the retained data.
@@ 2.3 LLM-driven pre-training corpus rewriting
Efforts to enhance code datasets using LLMs have gained traction. Austin et al. [12] explored LLM capabilities for program synthesis, finding that human feedback significantly improves code accuracy, but their work focused on synthesis rather than dataset rewriting. Cosmopedia [13] demonstrated the potential of synthetic data generation, using Mixtral-8×7B-Instruct [14] to create high-quality text corpora, though it did not address code-specific challenges.
Jain et al. [15] applied LLM-driven code transformations to instruction-tuning exemplars, focusing on variable renaming, modularization, and comment addition. While effective for fine-tuning, their approach is limited in scope and context compared to our SwallowCode pipeline. Their transformations address only a subset of stylistic improvements, whereas our Style-Guided Code Rewriting (SGCR) pipeline comprehensively enforces ten criteria from the Google Python Style Guide, including descriptive variable naming, effective type annotations, modular function design, error handling, and readability-focused formatting (see Section 3.3.1). Furthermore, our Self-Contained Optimization Rewriting (SCOR) pipeline introduces semantic enhancements—ensuring self-containment by resolving dependencies, optimizing algorithms, and transforming trivial snippets into instructive examples—which are absent in Jain et al.’s work (see Section 3.3.2). Additionally, Jain et al. target instruction-tuning data, a smaller and more curated dataset, whereas SwallowCode systematically rewrites large-scale pre-training corpora, a more challenging task due to the diversity and volume of code samples. Our pipeline integrates rigorous preprocessing with syntax error and linter-based filtering to ensure high-quality inputs for rewriting.
SwallowCode’s transform-and-retain paradigm upgrades low-quality code rather than discarding it, addressing limitations of existing datasets like The Stack v1 [1] and v2 [2], and Stack-Edu [3], which rely on filtering. By combining filtering, SGCR, and SCOR, SwallowCode produces a high-quality corpus that significantly improves performance on HumanEval and HumanEval+ benchmarks, advancing the state-of-the-art in code dataset curation.
2.4 Synthetic Data Generation for Code
Recent approaches, such as Magpie [16], leverage high-performing LLMs to generate synthetic instruction-tuning datasets from prompts and desired characteristics without relying on existing data. However, we opted against generating synthetic code data from scratch for SwallowCode due to two key limitations. First, previous work [17] demonstrates that low-diversity synthetic data restricts LLM performance. Second, achieving high diversity in synthetic code datasets requires diverse topics and keywords as seeds, as seen in Nemotron-4 340B’s synthetic instruction data [18]. For code, defining such seeds (e.g., varied algorithmic paradigms or problem domains) remains an open challenge, as no established methodology ensures sufficient diversity across programming tasks. Instead, our approach leverages high-quality code from The-Stack-v2, filtered for syntactic and stylistic rigor (Section 3.2), and applies LLM-driven rewriting to enhance quality while preserving the inherent diversity of real-world code. This transform-and-retain strategy maximizes data utility and avoids the risks of synthetic data homogeneity.
@@ 3 Construction of the code corpus
The development of SwallowCode is driven by an empirical and exploratory approach, informed by data ablation experiments evaluating each stage of the data processing pipeline, as illustrated in Figure 2. Specifically, we continually pre-trained the baseline model with the dataset only differing code text subset: before and after applying specific stage in the pipeline, and decide if we should adopt or discard the stage based on the evaluation results. In this section, we present the experimental results and describe the design choices shaping the pipeline. Please refer to Appendix H.1 for detailed results. To ensure full reproducibility, all code, configurations, and supporting materials are publicly available at https://github.com/rioyokotalab/swallow-code-math.
@@ 3.1 Experimental setup
To evaluate the impact of each design choice in our data processing pipeline, we conduct systematic data ablation studies. Each ablation trains a model that differs only in the target pre-training dataset, holding all other factors constant, including model architecture, parameter count, non-target data, total token budget, and hyperparameters. Specifically, we perform continual pre-training starting from Llama-3.1-8B, using a total of approximately 50 billion tokens, a maximum sequence length of 8,192, a global batch size of approximately 4 million tokens, and the Llama-3 tokenizer. The target dataset is processed with less than one epoch within each ablation. We evaluate model checkpoints approximately every 10 billion tokens using ten downstream benchmarks.
Continual pre-training is conducted using Megatron-LM (version core_r0.9.0) [19]. For evaluation, we utilize the BigCodeBench [20] and lm-evaluation-harness [21] on the following benchmarks: OpenBookQA [22], TriviaQA [23], HellaSwag [24], SQuAD 2.0 [25], XWINO [26], MMLU [27], GSM8K [28], BBH [29], HumanEval [30], and HumanEval+ [31]. The effectiveness of the code corpus is specifically assessed using HumanEval and HumanEval+. The pre-training data mixture consists of 84% multilingual text and 16% code. Detailed proportions and data source are provided in Appendix A.4.1, with comprehensive training hyperparameters reported in Appendix A.1. All ablation models, associated checkpoints, and supporting materials are publicly available in our Hugging Face repository to ensure full reproducibility.
@@ 3.2 Filtering
Figure 3: Performance comparison of filtering methods on HumanEval (left) and HumanEval+ (right) benchmarks. The original dataset (Stack v2) is compared against datasets processed with syntax error filtering, linter-based filtering, and LLM-based scoring (evaluated in ablation studies but not adopted). Syntax and linter-based filtering enhance code generation performance, while LLM-based scoring provides marginal gains, considering the computational cost.
@@ 3.2.1 Syntax error filtering
Despite the heuristic filtering in the BigCode project, the-stack-v2-train-smol-ids [2] includes Python code samples with invalid syntax according to Python 3.10 specifications. To address this, we apply syntax error filtering by compiling each code sample using Python’s built-in compile() function, discarding any samples that fail to compile. This process reduces the dataset from approximately 41 million to 37 million samples, a 9.7% reduction. As shown in Figure 3, removing syntactically invalid samples improves performance on HumanEval and HumanEval+  [^1] The performance trajectory of The Stack v2 exhibits an initial decline followed by a recovery, consistent with forgetting and subsequent adaptation observed in prior continual pretraining studies [32]. This behavior is not particularly noteworthy.. Consequently, we adopt syntax-error filtering as a standard step in all subsequent experiments.
@@ 3.2.3 LLM-based scoring filtering
Given these limited gains, we compare LLM-based scoring to our LLM-driven rewriting pipeline, which refines code snippets by enhancing clarity and correctness (Section 3.3). Comparing Figure 3 and Figure 4, although the rewriting pipeline requires 1.22 times the computational resources of LLM-based scoring (a 22% increase), it achieves significantly greater performance gains on HumanEval and HumanEval+ (detailed in Appendix F). Consequently, we do not incorporate LLM-based scoring in subsequent experiments, which favors the more effective and cost-efficient rewriting approach.
@@ 3.3 LLM-driven rewriting
Figure 4: Performance comparison of LLM-driven rewriting steps on HumanEval (left) and HumanEval+ (right) benchmarks. The pre-rewriting (syntax-error and linter-based filtering) is compared against SGCR and SCOR. SGCR improves performance by over 9 points, while SCOR, applied after SGCR, further enhances scores by over 5 points, demonstrating the effectiveness of stylistic and semantic rewrites in the SwallowCode pipeline.
Recent studies highlight the potential of large language models (LLMs) to transform training data, enhancing the performance of instruction-tuned models. Jain et al. [15] proposed three transformations—variable renaming, modularization, and plan annotation—for code cleaning, demonstrating their effectiveness for instruction tuning. However, their approach is limited to curated instruction-tuning data and a narrow set of stylistic improvements, lacking semantic optimizations or preprocessing integration.
To address these limitations, we developed the SwallowCode pipeline, which systematically rewrites large-scale pre-training corpora using two complementary LLM-driven processes: Style-Guided Code Rewriting (SGCR) and Self-Contained Optimization Rewriting (SCOR). SGCR revises Python code snippets based on the Google Python Style Guide  [^2] https://google.github.io/styleguide/pyguide.html, enforcing stylistic improvements like clear naming and modular design (Section 3.3.1). SCOR extends SGCR by ensuring self-containment and applying semantic optimizations, such as efficient algorithms and instructive examples (Section 3.3.2). To clarify, SwallowCode retains the code-only format of The-stack-v2-train-smol-ids [2], consisting solely of Python snippets without the text-code pair structure typical of instruction-tuning datasets. The rewriting pipeline, powered by Llama-3.3-70B-Instruct, enhances code quality through stylistic and semantic transformations without introducing instructional prompts or responses. Thus, performance gains on HumanEval and HumanEval+ (Section 3.4) stem from improved data quality, not from distillation of instruction tuning capabilities.
To illustrate the scope of these transformations, Table 1 compares the coverage of SGCR and SCOR against the approach of Jain et al. [15]. While SGCR addresses stylistic criteria like type hints, error handling, and docstrings, SCOR introduces semantic enhancements, including self-containment and optimization (algorithm and data structure), broadening the scope of the SwallowCode pipeline.
Table 1: Comparison of code transformation coverage for Jain et al. [15], SGCR, and SCOR.